# Calibrated and uncertain? Evaluating uncertainty estimates in binary classification models

### v1

@@ 2.4 Approximate Bayesian inference
If we consider $\mathbf{\nu}$ to be an analytical function $\nu(\mathbf{x})$ of $\mathbf{x}$, we can integrate over every possible function $\nu(\mathbf{x})$ to find the mean of the posterior
$\mathbb{E}[\nu(\mathbf{x})]=\int\nu(\mathbf{x})p(\mathbf{\nu(\mathbf{x})}|\mathcal{D}_{N})d\nu(\mathbf{x})=\int\nu(\mathbf{x})\frac{p(\mathcal{D}_{N}|\nu(\mathbf{x}))p(\nu(\mathbf{x}))}{p(\mathcal{D}_{N})}d\nu(\mathbf{x}),$
@@
At this point one must make some assumptions about what every possible $\nu$ means, which prior probabilities to assign to $\nu(\mathbf{x})$, and if necessary how to approximate this integral. The most common approach is to use parametric methods, which reduce the function-space by making it finite dimensional. On the other hand, nonparametric Bayesian inference methods allow us to keep the integral infinite dimensional. These latter methods are the “gold standard” for uncertainty estimation, as they are theoretically well founded and have properties that tend to align with how we wish uncertainty to behave, such as increased uncertainty in low density regions of data space. Examples of such methods are Gaussian Processes and Dirichlet Process mixture models, which make the infinite integral tractable by making assumptions which allow for using the nice analytical properties of Gaussian and Dirichlet distributions respectively.
Traditional parametric methods assume that the data is generated by simple well-known probability distributions with just a few parameters, and inference aims to determine the value of these parameters. These methods are extremely vulnerable to model misspecification. A more flexible and robust approach to parametric inference which supports a more flexible assumption on underlying distribution is to use Bayesian approximation methods such as variational inference. In variational inference, we assume $\nu(\mathbf{x})$ to be a member of a broad family of distributions $Q(\mathbf{c}|\mathbf{x},\boldsymbol{\theta})$ represented by some high-dimensional parametrization-vector $\boldsymbol{\theta}$ with a prior distribution $p(\boldsymbol{\theta})$, and assume that these capture most of the important variation across possible frequency distributions (Blei et al., 2016). We can for example use a neural network with weights and biases $\boldsymbol{\theta}$ and a normalized last layer as our family of distributions $Q(\mathbf{c}|\mathbf{x},\boldsymbol{\theta})$. It is then possible to incorporate the information from the training dataset $\mathcal{D}_{\text{train}}=(\mathcal{C}_{\text{train}},\mathcal{X}_{\text{train}})$ of size $N_{\text{train}}$ to calculate the posterior probability
@@
For high dimensional $\boldsymbol{\theta}$ or analytically intractable integrals, it is often convenient to numerically estimate the resulting integral by a crude Monte Carlo (MC) estimator with $\{\mathbf{\theta_{i}}\},i=1,2,..,T,$ sampled from its posterior distribution $p(\boldsymbol{\theta}|\mathcal{D_{\text{train}}})$, resulting in the approximation
$P(\mathbf{c}|\mathbf{x},\mathcal{D_{\text{train}}})=\int Q(\mathbf{c}|\mathbf{x},\boldsymbol{\theta})p(\boldsymbol{\theta}|\mathcal{D_{\text{train}}})d\boldsymbol{\theta}=\mathbb{E}_{\theta}[Q(\mathbf{c}|\mathbf{x},\boldsymbol{\theta})]$ (4)
@@
This quantity measures the spread of the estimated posterior, and it’s square root, $u_{Q}$, is what we use to quantify class probability uncertainty in this study. As the inference procedure should make the posterior converge towards the LRFD, the uncertainty should in general decrease with larger datasets, and may approach zero.
@@ 3 Methodology
In this study, we have created synthetic data representative of a generic two dimensional binary classification problem. We have specifically focused on a task with continuous, non-Gaussian, non-periodic unbounded underlying distributions of the features with significant overlap between classes. The simplicity of the data is intentional, aiming to facilitate an in-depth examination of the algorithms rather than the data itself. Toy models are highly idealized and extremely simplified models which are a frequently used but perhaps underappreciated tool employed by natural scientists (Reutlinger et al., 2018). While toy models may not always be very useful for solving specific complex problems, they provide an excellent basis for testing basic assumptions and increase understanding. They are frequently used in statistical modeling studies but are less commonly seen in modern deep learning literature. It should be noted that the classification task is not trivial even if it is low dimensional. The class probability depends on the input features in a geometrically symmetric way, which none of the models can easily represent. The primary motivation for using toy data in this study is that by knowing the precise expression for the joint probability and the associated class probabilities as functions of the data features, we can exactly evaluate the estimates produced by the machine learning algorithms.
We investigated six different ML classification algorithms with probability and uncertainty estimates: neural network ensemble (NNE), neural network ensemble with conflictual loss (CL) (Fellaji et al., 2024), neural network using evidential deep learning (EDL) (Sensoy et al., 2018), neural network with Monte Carlo Dropout (MCD) (Gal and Ghahramani, 2016), Gaussian Process classification (GP)(Rasmussen and Williams, 2006) and a Dirichlet Process Mixture Model (DPMM) (Li et al., 2019). Four are deep learning algorithms while the other two are nonparametric Bayesian inference models. Details of the class probability and uncertainty estimation as well as extensive hyperparameter sweeps for each algorithm can be found in Appendices A and B.
@@ 3.1 Research questions and evaluation metrics
Our first research question is if estimated probabilities are calibrated. This has been studied thoroughly and different metrics of calibration such as expected calibration error (Naeini et al., 2015; Guo et al., 2017; Minderer et al., 2021) have been devised. However, the majority of these studies are done on data where the LRFD is not known and must be approximated using binning schemes on finite test data. In our case, Q1 can be investigated by comparing the probability estimates for our test set (by default in-distribution data) with the LRFD of the data. Because of the symmetries of a binary classification problem where $P(c^{1}=1|\mathbf{x})=1-P(c^{2}=1|\mathbf{x})$ we will only show the results for class 2. The differences are evaluated by plotting the estimates next to the LRFD and by calculating five summary statistics of calibration on the test set, specifically
@@
The second research question aims to answer if the uncertainties produced by the algorithms actually decrease when we increase the number of data points. In the study of Fellaji et al. (2024), where ensembles with conflictual loss are introduced, the authors call this the data-related principle of epistemic uncertainty and find NNE, MCD and EDL to not follow this principle, which is the motivation for their introduction of the CL algorithm. To evaluate if this is the case for our data, we use the average of the uncertainty estimates over the whole test set as a metric. Because of the stochastic nature of both data and training, we expect this metric to fluctuate, but it should in general go down as we add more data to the machine learning models.
The last research question is related to generalization and the ability of flexible models to make sensible predictions outside the training distribution. Ideally, we would want our uncertainty estimates to be high for data which is very different from the data previously seen by the learning algorithm (Hein et al., 2019). In our case, this means that we tested to see if probability estimates approached 0.5 and uncertainty was high in the tails of our distribution. We evaluate this by calculating the average probability estimates and uncertainty on an extrapolation test set, as well as looking at the behavior of estimates in the tails of the in-distribution test set.
@@ 4 Dataset and experiment
Figure 1: The conditional and marginal distributions for the two datasets. The upper row shows the distribution from which we sampled dataset A, with parameters $\alpha_{1},\theta_{1}=[2,5]$, $\alpha_{2},\theta_{2}=[6,3]$. The lower row shows the distribution from which we sampled dataset B, with parameters $\alpha_{1},\theta_{1}=[2,3]$ and $\alpha_{2},\theta_{2}=[4,3]$. The panels from left to right show the conditional distributions $P(c|r)$, $p(r|c)$, and the marginal distributions $p(r)$ and $P(c)$.
@@
Figure 2: Subsets of the training dataset with $N_{train}=250,5000$ and 1000010000 data points. The upper row shows the data points of class 1 in red and the lower row shows the data points of class 2 in blue. The left panel shows dataset A and the right panel shows dataset B.
@@
Each data point $d$ consists of features $\mathbf{c}\in\{[1,0],[0,1]\}$, and $\mathbf{x}=[x^{1},x^{2}]\in\mathbf{R}^{2}$ with polar coordinates $r=\sqrt{(x^{1})^{2}+(x^{2})^{2}}$ and $\phi=\arctan(x^{2}/x^{1})$. To facilitate visual interpretation we have chosen a distribution with radial symmetry. $\mathbf{x}$ is distributed such that $p(r|c^{j}=1)=\Gamma(\alpha_{j},\eta_{j},r)$, where $\Gamma$ is the gamma distribution with shape parameter $\alpha$ and scale parameter $\eta$. The angular density $p(\phi|c^{j}=1)=1/2\pi$ for both classes. Two different classification datasets, dataset A and dataset B, were generated by setting the parameters of the gamma distributions to different values. For dataset A the parameter values are $\alpha_{1},\eta_{1}=[2,5]$, $\alpha_{2},\eta_{2}=[6,3]$ where the index indicates the class, and for dataset B the parameter values are $\alpha_{1},\eta_{1}=[2,3]$, $\alpha_{2},\eta_{2}=[4,3]$. The marginal class probabilities $P(\mathbf{c})$ are uniform for both datasets. The distributions of the data are shown in Figure 1. These specific distributions were chosen because of the different behavior for $P(\mathbf{c}|r)$ in the beginning of the tails, around $r=30$. Train, validation and test sets with 10000, 5000, and 10000 samples, respectively, were generated. To check estimated probabilities and uncertainties for OOD test points, a grid in polar coordinates was created by letting $\phi_{i}\in\{0,\frac{2}{5}\pi,\frac{4}{5}\pi,\frac{6}{5}\pi,\frac{8}{5}\pi\}$ and $r_{i}$ be 26 equally-spaced values in log-space between 700 and 1000 for each polar angle, resulting in 130 data points.
@@
Each model was trained on the same subsets of increasing amounts of training data with $N_{\text{train}}=$ 250, 500, 1000, 2000, 3000, 5000 and 10000. Figure 2 shows the training data for 250, 1000 and 10000 training points. As the gradient descent optimization algorithm is sensitive to initialization values of its parameters and the order in which training data is seen, 20 separate training runs were performed for each NN algorithm and the one with the lowest cross entropy loss on the validation set was kept. The trained models were then used to estimate the class probability and uncertainty on the test set and the extrapolation grid. Details on training and implementation can be found in Appendix A.
@@ 5 Results
Figure 3: Estimated probabilities (top row) and uncertainties (bottom row) for class 2 for the different algorithms for dataset A as a function of radius $|\mathbf{x}|$. The error bars indicate the entire spread of the data over polar angle $\phi$. The long-run frequency distribution (solid black line) is plotted for reference.
@@
Figure 4: Estimated probabilities (top row) and uncertainties (bottom row) for class 2 for the different algorithms for dataset B as a function of radius $|\mathbf{x}|$. The error bars indicate the entire spread of the data over polar angle $\phi$. The long-run frequency distribution (solid black line) is plotted for reference.
@@
5.1 Calibration results

### v2

@@ 2.5 Approximate Bayesian inference
Going back to the expression for the predictive posterior in Equation 2, we will consider $\mathbf{\nu}$ to be an analytical function $\nu(\mathbf{x})\in\mathcal{F}$, where $\mathcal{F}$ is the set containing all possible functions $\mathbf{\nu}$. We can then integrate over the in-principle infinite dimensional function-space $\mathcal{F}$ to find the mean of the posterior at any point $\mathbf{x}$
$\mathbb{E}[\nu(\mathbf{x})]=\int_{\mathcal{F}}\nu(\mathbf{x})p(\mathbf{\nu(\mathbf{x})}|\mathcal{D}_{N})d\nu(\mathbf{x})=\int_{\mathcal{F}}\nu(\mathbf{x})\frac{p(\mathcal{D}_{N}|\nu(\mathbf{x}))p(\nu(\mathbf{x}))}{p(\mathcal{D}_{N})}d\nu(\mathbf{x}),$ (6)
@@
At this point one we can implement any problem-specific assumptions and prior knowledge to define what every possible $\nu$ means, which prior probabilities to assign to $\nu(\mathbf{x})$, and if necessary how to approximate this integral. In scientific classification problems with real-valued input features $\mathbf{x}$, it is common to assume that $P(c^{j}=1|\mathbf{x})$ as a function of $\mathbf{x}$ is continuous and smooth. This is equivalent to making the assumption that the densities $p(\mathbf{x}|c^{j}=1)$ are continuous and smooth, as $P(c^{j}=1|\mathbf{x})\propto p(\mathbf{x}|c^{j}=1)$.
The most common approach to inferences over function-space is to use parametric methods, which reduce the infinite-dimensional function-space by making it finite dimensional. On the other hand, nonparametric Bayesian inference methods allow us to keep the integral infinite dimensional. These latter methods are often seen as the “gold standard” for uncertainty estimation, as they are theoretically well founded and have properties that tend to align with how we wish uncertainty to behave, such as increased uncertainty in low density regions of data space. However, implementing high dimensional nonparametric inference models is notoriously difficult and the resulting algorithms often need computationally expensive Monte Carlo simulations which make them less useful in practice. Examples of such methods are Gaussian Processes and Dirichlet Process mixture models, which make the infinite integral tractable by making assumptions which allow for using the nice analytical properties of Gaussian and Dirichlet distributions respectively.
Traditional parametric methods assume that the data is generated by simple well-known probability distributions with just a few parameters, and inference aims to determine the value of these parameters. These methods are extremely vulnerable to model misspecification. An alternative approach to classical parametric inference is to use approximate Bayesian nonparametric inference methods such as Bayesian neural networks. In approximate Bayesian inference over function spaces, we let the functions $\nu(\mathbf{x})$ in Equation 6 be the members of a broad family of distributions $Q(\mathbf{c}|\mathbf{x},\boldsymbol{\theta})$ represented by some high-dimensional parametrization-vector $\boldsymbol{\theta}$ with a prior distribution $p(\boldsymbol{\theta})$. We assume that this family of distributions capture most of the important variation across possible frequency distributions of the posterior (Blei et al., 2016). We can for example use a neural network with weights and biases $\boldsymbol{\theta}$ and a normalized last layer as our family of distributions $Q(\mathbf{c}|\mathbf{x},\boldsymbol{\theta})$. It is then possible to incorporate the information from the training dataset $\mathcal{D}_{\text{train}}=(\mathcal{C}_{\text{train}},\mathcal{X}_{\text{train}})$ of size $N_{\text{train}}$ to calculate the posterior probability
@@
For high dimensional $\boldsymbol{\theta}$ or analytically intractable integrals we have to estimate this integral, for example by a Monte Carlo (MC) estimate with $\{\mathbf{\theta_{i}}\},i=1,2,..,T,$ sampled from its posterior distribution $p(\boldsymbol{\theta}|\mathcal{D_{\text{train}}})$. This results in the approximation
$P(\mathbf{c}|\mathbf{x},\mathcal{D_{\text{train}}})\approx\int Q(\mathbf{c}|\mathbf{x},\boldsymbol{\theta})p(\boldsymbol{\theta}|\mathcal{D_{\text{train}}})d\boldsymbol{\theta}=\mathbb{E}_{\theta}[Q(\mathbf{c}|\mathbf{x},\boldsymbol{\theta})]$ (7)
@@
This quantity measures the spread of the estimated posterior, and its square root, $u_{Q}$, is what we use to quantify class probability uncertainty in this study. As the inference procedure should, in the case of no approximations, make the posterior converge towards the LRFD, the uncertainty should in general decrease with larger datasets, and may approach zero. If we do not see this behavior in our model we should suspect that the model is misspecified, our approximations are too dominant, or there is some error in the implementation of the algorithm.
@@ 3 Methodology
In this study, we have created synthetic data representative of a generic two dimensional binary classification problem. We have specifically focused on a task with continuous, non-Gaussian, non-periodic unbounded underlying distributions of the features with significant overlap between classes. The simplicity of the data is intentional, aiming to facilitate an in-depth examination of the algorithms rather than the data itself. Toy models are highly idealized and extremely simplified models which are a frequently used but perhaps underappreciated tool employed by natural scientists (Reutlinger et al., 2018). While toy models may not always be very useful for solving specific complex problems, they provide an excellent basis for testing basic assumptions and increase understanding. They are frequently used in statistical modeling studies but are less commonly seen in modern deep learning literature. It should be noted that the classification task is not trivial even if it is low dimensional. The class probability depends on the input features in a geometrically symmetric way, which none of the models can easily represent. The primary motivation for using toy data in this study is that by knowing the precise expression for the asymptotic joint probability distribution and the associated conditional class probabilities (referred to as the LRFD, $\nu_{\infty}$, in this paper) as functions of the data features, we can exactly evaluate the estimates produced by the machine learning algorithms.
We investigated six different ML classification algorithms with probability and uncertainty estimates: neural network ensemble (NNE), neural network ensemble with conflictual loss (CL) (Fellaji et al., 2024), neural network using evidential deep learning (EDL) (Sensoy et al., 2018), neural network with Monte Carlo Dropout (MCD) (Gal and Ghahramani, 2016), Gaussian Process classification (GP)(Rasmussen and Williams, 2006) and a Dirichlet Process Mixture Model (DPMM) (Li et al., 2019). Four are deep learning algorithms while the other two are nonparametric Bayesian inference models. There exists many other algorithms for uncertainty quantification in deep learning. Some of the more well known are Bayesian neural networks with posteriors calculated using MCMC, Laplace approximations, variational inference, SWA-Gaussian and conformal predictions. It is not the goal of this paper to test all of them, but rather to create a framework for testing these kind of algorithms and evaluate a limited but representative sample of popular methods for uncertainty quantification in deep learning. Details of the class probability and uncertainty estimation as well as extensive hyperparameter sweeps for each algorithm can be found in Appendices A and B.
@@ 3.1 Research questions and evaluation metrics
Our first research question is if estimated probabilities are calibrated. This topic has been studied thoroughly and different metrics of calibration such as expected calibration error (Naeini et al., 2015; Guo et al., 2017; Minderer et al., 2021) have been devised to evaluate calibration properties of classifiers. However, the majority of these studies are performed on real data where the LRFD is not known and must be approximated using binning schemes on finite test data as described in Section 2.3. In our case, Q1 can be investigated by comparing the probability estimates for our test set (by default in-distribution data) with the LRFD of the data. Because of the symmetries of a binary classification problem where $P(c^{1}=1|\mathbf{x})=1-P(c^{2}=1|\mathbf{x})$ we will only show the results for class 1. The differences are evaluated by plotting the estimates next to the LRFD and by calculating five summary statistics of calibration on the test set, specifically
@@
In Appendix C we also present reliability diagrams as an extra measure of (marginal) calibration.
The second research question aims to answer if the uncertainties produced by the algorithms actually decrease when we increase the number of data points. In the study of Fellaji et al. (2024), where ensembles with conflictual loss are introduced, the authors call this the data-related principle of epistemic uncertainty and find NNE, MCD and EDL to not follow this principle, which is the motivation for their introduction of the CL algorithm. To evaluate if this is the case for our data, we use the average of the uncertainty estimates calculated over the test set as a metric. Because of the stochastic nature of both data and training, we expect this metric to fluctuate, but it should in general go down as we add more data to the machine learning models.
The last research question is related to generalization properties and the ability of flexible models to make sensible predictions outside the training distribution. Ideally, we would want our uncertainty estimates to be high for data which is very different from the data previously seen by the learning algorithm (Hein et al., 2019). In our case, this means that we tested to see if probability estimates approached 0.5 and uncertainty was high in the tails of the data distributions. In scientific applications it can be useful to be able to detect both that a new data point is located in the start of the tail of the distribution as well as very far away. This is especially important in safety-critical applications where both small and big shifts from what we expect to see can be problematic. We evaluate estimates for both types of OOD data by calculating the average probability estimates and uncertainty for two different extrapolation test set, one relatively near the peak of the data density and one located an order of magnitude further away. We also looked at the behavior of the same estimates in the tails of the in-distribution test set.
@@ 4 Dataset and experiment
Figure 1: The conditional and marginal generating distributions for the two datasets. The upper row shows the distribution from which we sampled dataset A, with parameters $\alpha_{1},\eta_{1}=[5,2]$, $\alpha_{2},\eta_{2}=[3,6]$. The lower row shows the distribution from which we sampled dataset B, with parameters $\alpha_{1},\eta_{1}=[3,2]$ and $\alpha_{2},\eta_{2}=[3,4]$. The panels from left to right show the conditional distributions $P(c^{j}=1|r,\alpha_{1},\alpha_{2},\eta_{1},\eta_{2},P_{c})=\nu^{j}_{\infty}(r)$, $p(r|c,\alpha_{j},\eta_{j})$, and the marginal distributions $p(r|\alpha_{1},\alpha_{2},\eta_{1},\eta_{2},P_{c})$ and $P(c)$.
@@
Figure 2: Subsets of training datasets A (left) and B (right) with $N_{train}=250,5000$ and 1000010000 data points. The upper row shows the data points of class 1 in red and the lower row shows the data points of class 2 in blue.
@@
Each data point $d$ consists of features $\mathbf{c}\in\{[1,0],[0,1]\}$, and $\mathbf{x}=[x^{1},x^{2}]\in\mathbb{R}^{2}$ with polar coordinates $r=\sqrt{(x^{1})^{2}+(x^{2})^{2}}$ and $\phi=\arctan(x^{2}/x^{1})$. To facilitate visual interpretation we have chosen a generating distribution with radial symmetry. $\mathbf{x}$ is distributed such that $p(r|c^{j}=1)=\Gamma(r|\alpha_{j},\eta_{j})$, where $\Gamma$ is the gamma distribution with shape parameter $\alpha_{j}$ and scale parameter $\eta_{j}$. The angular density is a uniform density given by $p(\phi|c^{j}=1)=p_{\phi}=1/2\pi$ for both classes. Two different classification datasets, dataset A and dataset B, were generated by setting the parameters of the gamma distributions to different values. For dataset A the parameter values are $\alpha_{1},\eta_{1}=[5,2]$, $\alpha_{2},\eta_{2}=[3,6]$ where the index indicates the class, and for dataset B the parameter values are $\alpha_{1},\eta_{1}=[3,2]$, $\alpha_{2},\eta_{2}=[3,4]$. The marginal class probabilities $P(\mathbf{c})=P_{c}=0.5$ are uniform for both datasets. The generating distributions of the data are shown in Figure 1. These specific distributions were chosen because of the different behavior for $P(\mathbf{c}|r)$ in the beginning of the tails, around $r=30$ and $r=20$ for dataset A and B respectively. Train, validation and test sets with 10000, 5000, and 10000 samples, respectively, were generated by random sampling from the joint distribution $p(\mathbf{c},x^{1},x^{2}|\alpha_{1},\alpha_{2},\eta_{1},\eta_{2},p_{\phi},P_{c})$ by first sampling the class labels, then the angles and finally the radial positions of the data points [^3] The datasets and code used for this study is made available at https://github.com/choisant/foundational-modeling..
To check estimated probabilities and uncertainties for OOD test points very far from the density peak of the data, a grid in polar coordinates was created by letting $\phi_{i}\in\{0,\frac{2}{5}\pi,\frac{4}{5}\pi,\frac{6}{5}\pi,\frac{8}{5}\pi\}$ and $r_{i}$ be 26 equally-spaced values in log-space for $700\leq r\leq 1000$ for each polar angle, resulting in 130 data points. The same procedure was used to create a test set for OOD data points that were relatively close to the data, specifically in the range $80\leq r\leq 120$. In the text we will refer to these two OOD test sets as ”OOD-near” and ”OOD-far”.
@@
Each model was trained on the same subsets of increasing amounts of training data with $N_{\text{train}}=$ 250, 500, 1000, 2000, 3000, 5000 and 10000. Figure 2 shows the training data with 250, 1000 and 10000 training points. As the gradient descent optimization algorithm is sensitive to initialization values of its parameters and the order in which training data is seen, 20 separate training runs were performed for each NN algorithm and the one with the lowest cross entropy loss on the validation set was kept. The main results for the deep learning models are obtained using a fully connected neural network architecture and the commonly used ReLU activation function. The trained models were then used to estimate the class probability and uncertainty on the test set and the extrapolation grid. Details on training and implementation can be found in Appendix A. An extended study over varying hyperparameters and architecture choices are presented in Appendix B and C.
@@ 5 Results
Figure 3: Estimated probabilities (top row) and uncertainties (bottom row) for class 1 for the different algorithms for dataset A as a function of radius $|\mathbf{x}|$. The error bars indicate the entire spread of the data over polar angle $\phi$, while the markers indicate the sample average. The long-run frequency distribution (solid black line) is plotted for reference.
@@
Figure 4: Estimated probabilities (top row) and uncertainties (bottom row) for class 1 for the different algorithms for dataset B as a function of radius $|\mathbf{x}|$. The error bars indicate the entire spread of the data over polar angle $\phi$. The long-run frequency distribution (solid black line) is plotted for reference.
@@
(a) Dataset A
(b) Dataset B
Figure 6: Calibration metrics of the test set as a function of number of training data points $N_{\text{train}}$ (as in Figure 5), but calculated for the tails ($|\mathbf{x}|>30$ for dataset A and $|\mathbf{x}|>20$ for dataset B) of the test sets, as indicated by the dashed black line in Figure 2.
5.1 Q1: Calibration
First we attempt to answer Q1 by evaluating if the estimated probabilities are calibrated. We find that the probability estimates for all the different models are reasonably well calibrated with respect to input features for the majority of test data points, with calibration getting better for larger values of $N_{\text{train}}$. In the tails of the test data distribution, the algorithms are not well calibrated unless the LRFD aligns with the extrapolation behavior of the algorithm. Global calibration metrics and reliability diagrams support this statement quantitatively.